EDITORIAL

Terry Goodman

We are announcing that the Missouri Journal of Mathematical Sciences will be accepting well-motivated and sharply focused short notes. These are often gems that provide a new proof of an old theorem, a novel presentation of a familiar theme, or a lively discussion of a single issue. Notes may convey new mathematical insights and effective teaching strategies for college mathematics instruction. Papers should be submitted electronically as a single file to either Shing So or Curtis Cooper. You can find their email addresses on the inside front cover of this issue. It is strongly recommended that authors prepare their papers with AMS-TEX, LATEX, or TEX and files for figures in EPS format. Manuscripts should contain a short abstract, 2000 Mathematics Subject Classification, key words, author’s name, address, email address, and affiliation. Joint authors should indicate on the manuscript which author should receive the proof. Manuscripts should be typewritten in a double-spaced, wide margin format to allow for editorial remarks and should not exceed 6 pages of text. We look forward to publishing this new section in MJMS.

By now you have seen our new cover design. We would like to acknowledge and thank Hang Chen for his time and hard work in creating the new MJMS cover. We appreciate his artistic efforts to enhance the appearance of the MJMS.

Finally, as is our custom, we want to acknowledge those who have served as referees for MJMS for the past year. Their contribution to the journal is so very important and we appreciate the time and energy they give to their task of reviewing submitted manuscripts.

List of Referees

Om Prakash Agrawal, Southern Illinois University Ibrahim Assem, University of Sherbrooke, Canada Nick Baeth, University of Central Missouri Fabio Di Benedetto, Genova University, Italia James Benedict, Augusta State College Renato Betti, Politecnico di Milano, Italia Stephen Clark, University of Missouri-Rolla Kumer Das, Lamar University Joseph B. Dence, University of Missouri-St. Louis Roman Dwilewicz, University of Missouri-Rolla S. S. Dragomir, Victoria University, Australia Ahmad El-guindy, Texas A & M University

List of Referees (cont.)

Gail Fahoome, Wayne State University Marco Frontini, Politecnico di Milano, Italia Ghiocel Groza, Technical University of Civil Engineering, Romania John Hannah, University of Canterbury, New Zealand Joe Howard, New Mexico Highlands University James Howie, Heriot-Watt University, UK Eric Howington, Coastal Carolina University Dustin Jones, Sam Houston State University Charles Lanski, University of Southern California Vy Le, University of Missouri-Rolla Rhonda McKee, University of Central Missouri Alexander Moreto, Universitat de Valencia, Spain Ollie Nanyes, Bradley University Don Redmond, Southern Illinois University Sr. Barbara Reynolds, Cardinal Stritch University Ken Richardson, Texas Christian University Leanne D. Robertson, Smith College Robert Roe, University of Missouri-Rolla David Rose, Florida Southern University Philip Schultz, University of Western Australia Temba Shonhiwa, Wits University, South Africa Neil O. Smith, Augusta State College Ashish Srivastava, Ohio University Scott Thatcher, Truman State University Stefan Veldsman, Wits University, South Africa Haohao Wang, Southeast Missouri State University Lianwen Wang, University of Central Missouri Lawrence Washington, University of Maryland Kenneth Williams, Carleton University Mahmoud Yousef, University of Central Missouri